rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1217 — FS#5224 — tou59-1-3k

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
We detected a problem of optical on tou59-1-3k :

%SFF8472-5-THRESHOLD_VIOLATION: Te0/1: Rx power low alarm; Operating value: -16.0 dBm, Threshold value: -15.2 dBm.

We will intervene tomorrow morning at 10h00. An interruption of about 1 hour is predicted. The customers on the NRA of Roncq will be impacted during this maintenance.


Date:  Friday, 11 March 2011, 15:57PM
Reason for closing:  Done
Comment by OVH - Friday, 11 March 2011, 13:21PM

We will ask for the intervention. The NRA of roncq will be available during the tests and the optic replacement.


Comment by OVH - Friday, 11 March 2011, 13:24PM

We have just cleaned out the optic bloc in addition to the fibre, the optic power supply of reception is back to normal. The NRA of roncq is up again.

Optical Optical
Temperature Voltage Tx Power Rx Power
Port (Celsius) (Volts) (dBm) (dBm)
--------- ----------- ------- -------- --------
Te0/1 44.8 3.38 -1.6 -11.2